# Double Robustness of Local Projections

### v2

@@ 6 Conclusion
    Researchers who use LP should control for those lags of the data that are strong predictors of the outcome or impulse variables. This is important even if the researcher directly observes a near-perfect proxy for the shock of interest. However, it is not necessary to get the lag length exactly right to achieve correct coverage. To select the number of lags to control for in the LP, it suffices to run a VAR in all variables used in the analysis and select the lag length that minimizes conventional information criteria (such as AIC). Our results complement the finding of Montiel Olea and Plagborg-Møller (2021) that lag-augmented LP confidence intervals are also more robust than VAR intervals to persistence in the data and to the length of the impulse response horizon.
@@
    There is no free lunch for VARs: if an estimated VAR yields confidence intervals that are substantially narrower than the corresponding LP intervals, we recommend increasing the VAR lag length until that is no longer the case, to guarantee robust confidence interval coverage. Conventional tests of correct VAR specification do not suffice to guard against coverage distortions.
Is there a way forward for VAR inference, beyond just including a large number of lags? We showed how to construct a VAR confidence interval with a bias-aware critical value that robustly controls coverage, but found that it will typically lead to wider confidence intervals than LP. Another option would be to estimate VARMA models rather than pure VARs, though this would be computationally expensive, and the bias-variance trade-off relative to LPs is unclear. In principle, VAR procedures may work better under additional restrictions on the misspecification, such as shape restrictions on the impulse response functions. [^12] Given any convex parameter space for the misspecification MA polynomial $\alpha(L)$, the worst-case bias of the VAR estimator (see Proposition 3.2) can be computed using convex programming. However, it appears that detailed application-specific restrictions would be required to generate a negligible worst-case bias, since we have shown that the least favorable misspecification in our baseline analysis is economically plausible. Rather than restricting the parameter space, future research could instead investigate weakening the coverage requirement, e.g., only requiring a certain coverage probability on average over a set of horizons (Armstrong, Kolesár, and Plagborg-Møller, 2022), or by changing the target for inference from the true impulse response function to a smooth projection of this function (Genovese and Wasserman, 2008). Finally, a subjectivist Bayesian VAR modeler need only worry about our negative results if their prior on potential misspecification attaches significant weight to MA processes that imply large VAR biases.
@@ A.1 Least favorable misspecification
Figure A.1 plots some numerical examples of the least favorable MA polynomial $\alpha^{\dagger}(L;h,M)=\sum_{\ell=1}^{\infty}\alpha^{\dagger}_{\ell,h,M}L^{\ell}$ discussed in Section 4.2. We focus here on a univariate local-to-AR(1) model $y_{t}=\rho y_{t-1}+[1+T^{-1/2}\alpha(L)]\varepsilon_{t}$, though unreported numerical experiments suggest that the qualitative features mentioned below also apply to multivariate models. Recall that the least favorable MA coefficients depend on the horizon $h$ of interest, while $M$ only influences the overall scale of the coefficients, and not their shape as a function of $\ell$. The figure shows that the shape of the coefficients either takes the form of a hump or of a single zig-zag pattern, with the largest absolute value of the coefficients generally occurring at $\ell=h$. Notice that we can flip the signs of all coefficients without changing the bias.
@@ A.2 More efficient bias-aware confidence interval
Generalizing the bias-aware VAR confidence interval in Section 4.3, consider a bias-aware confidence interval that is centered at the model averaging estimator $\hat{\theta}_{h}(\omega)=\omega\hat{\beta}_{h}+(1-\omega)\hat{\delta}_{h}$ from Corollary 4.2:
@@
Impose Assumption 2.1, $\zeta=1/2$, and $\operatorname{aVar}(\hat{\delta}_{h})>0$. Then, for any $\omega\in[0,1]$,
@@
The result follows from Propositions 3.1, 3.2, A.2 and 4.1 and the same calculations as in the proof of Corollary 4.2. ∎
@@
Figure A.3 shows this optimal weight as a function of $M$ and the relative asymptotic standard deviation of the VAR and LP estimators, while Figure A.3 shows the length of the resulting optimal bias-aware confidence interval relative to the length of the conventional LP interval. We see that, for $M\geq 2$, there is little gain from reporting the optimal bias-aware interval rather than the LP interval, regardless of the relative precision of VAR and LP. An additional observation is that, for $M\geq 1.5$, the length-optimal $\omega^{*}$ is numerically close to the MSE-optimal weight $M^{2}/(1+M^{2})$ derived in Corollary 4.2.
Figure A.2: Length-optimal weight on LP in bias-aware confidence interval. Significance level $a=10\%$. Horizontal axis: relative asymptotic standard deviation of LP vs. VAR. Different lines: different bounds $M$ on $\|\alpha(L)\|$.
Figure A.3: Relative length of optimal bias-aware confidence interval vs. conventional LP interval. Significance level $a=10\%$. Horizontal axis: relative asymptotic standard deviation of LP vs. VAR. Different lines: different bounds $M$ on $\|\alpha(L)\|$.
@@ A.3 Covariance structure of LP and VAR estimators
The following result provides the asymptotic variance-covariance matrix of the LP and VAR estimators in the general multi-dimensional set-up of Section 4.4. Define $\Psi_{i^{*},j^{*},h}$ as in Proposition 3.2, but making the dependence on $(i^{*},j^{*})$ explicit in the notation.
@@
Impose Assumption 2.1, with part (iii) holding for all shock indices $j_{1}^{*},\dots,j_{k}^{*}$. Then for any $a,b\in\{1,\dots,k\}$,
@@
and the “$\operatorname{aCov}$” notation is understood to refer to elements of the asymptotic variance-covariance matrix in (4.2). In particular, $\operatorname{aCov}(\hat{\beta}_{i_{a}^{*},j_{a}^{*},h_{a}}-\hat{\delta}_{i_{a}^{*},j_{a}^{*},h_{a}},\hat{\delta}_{i_{b}^{*},j_{b}^{*},h_{b}})=0$.
@@ Appendix B Proofs
Lemmas whose name begins with “E” can be found in the Supplemental Appendix.
B.1 Proof of Proposition 3.1
@@
$y_{i^{*},t+h}=\theta_{h,T}\varepsilon_{j^{*},t}+\underline{B}_{h,y}^{\prime}\underline{y}_{j^{*},t}+B_{h,y}^{\prime}y_{t-1}+\xi_{h,i^{*},t}+T^{-\zeta}\Theta_{h}(L)\varepsilon_{t},$ (B.1)
where the precise expressions for the coefficient matrices are given in LABEL:thm:lp_coefficients_aux. The lemma also shows that the $1\times n$ two-sided lag polynomial $\Theta_{h}(L)=\sum_{\ell=-\infty}^{\infty}\Theta_{h,\ell}L^{\ell}$ is absolutely summable and satisfies $\Theta_{h,0,j^{*}}=0$. That is, $\Theta_{h}(L)\varepsilon_{t}$ is independent of $\varepsilon_{j^{*},t}$ (but not of $\varepsilon_{j^{*},s}$ for $s\neq t$).
Let $\hat{x}_{h,t}$ be the residual in a regression of $y_{j^{*},t}$ on $\underline{y}_{j^{*},t}$ and $y_{t-1}$. By definition, $\hat{x}_{h,t}$ is in-sample orthogonal to $\underline{y}_{j^{*},t}$ and $y_{t-1}$. Hence,
@@
$=\theta_{h,T}+\frac{T^{-1}\sum_{t=1}^{T-h}(\xi_{h,i^{*},t}+T^{-\zeta}\Theta_{h}(L)\varepsilon_{t})\varepsilon_{j^{*},t}+o_{p}(T^{-1/2})+O_{p}(T^{-2\zeta}+T^{-1/2-\zeta})}{\sigma_{j^{*}}^{2}+o_{p}(1)}+o_{p}(T^{-1/2})$
by LABEL:thm:x(LABEL:itm:thm_x_iv) and (LABEL:itm:thm_x_v)
$=\theta_{h,T}+\frac{T^{-1}\sum_{t=1}^{T-h}(\xi_{h,i^{*},t}+T^{-\zeta}\Theta_{h}(L)\varepsilon_{t})\varepsilon_{j^{*},t}}{\sigma_{j^{*}}^{2}+o_{p}(1)}+o_{p}(T^{-1/2}).$
Finally, LABEL:thm:lp_coefficients_aux shows that
$T^{-1}\sum_{t=1}^{T-h}(\Theta_{h}(L)\varepsilon_{t})\varepsilon_{j^{*},t}=O_{p}(T^{-1/2}).$
Thus, we conclude that
$T^{-1}\sum_{t=1}^{T-h}T^{-\zeta}(\Theta_{h}(L)\varepsilon_{t})\varepsilon_{j^{*},t}=O_{p}(T^{-1/2-\zeta})=o_{p}(T^{-1/2}),$
which completes the proof. ∎
B.2 Proof of Proposition 3.2
@@
LABEL:thm:A shows that $\hat{A}-A=O_{p}(T^{-\zeta}+T^{-1/2})$. Since it is known that
@@
see for example Magnus and Neudecker (2007, Table 7, p. 208), the delta method gives
@@
where $S$ was defined in Assumption 2.1. LABEL:thm:nu shows that
$\hat{\nu}-H_{\bullet,j^{*}}=\frac{1}{\sigma_{j^{*}}^{2}}T^{-1}\sum_{t=1}^{T}\xi_{0,t}\varepsilon_{j^{*},t}+o_{p}(T^{-1/2}).$
@@
B.3 Proof of Proposition 4.2
Define $\tilde{\alpha}_{\ell}=D^{-1/2}\alpha_{\ell}D^{1/2}$ for all $\ell\geq 1$. Notice that $\|\alpha(L)\|^{2}=\sum_{\ell=1}^{\infty}\|\tilde{\alpha}_{\ell}\|^{2}$.
By Proposition 3.2, we have $\operatorname{aBias}(\hat{\delta}_{i^{*},j^{*},h})=\sum_{\ell=1}^{\infty}\operatorname*{trace}(\Xi_{i^{*},j^{*},h,\ell}\tilde{\alpha}_{\ell})$, where
$\Xi_{i^{*},j^{*},h,\ell}\equiv D^{1/2}H^{\prime}(A^{\prime})^{\ell-1}S^{-1}\Psi_{i^{*},j^{*},h}HD^{1/2}-\mathbbm{1}(\ell\leq h)D^{-1/2}e_{j^{*},m}e_{i^{*},n}^{\prime}A^{h-\ell}HD^{1/2}.$
Since $\operatorname*{trace}(\Xi_{i^{*},j^{*},h,\ell}\tilde{\alpha}_{\ell})=\operatorname*{vec}(\Xi_{i^{*},j^{*},h,\ell})^{\prime}\operatorname*{vec}(\tilde{\alpha}_{\ell}^{\prime})$, we can write
$\operatorname{aBias}(\hat{\boldsymbol{\delta}})=\sum_{\ell=1}^{\infty}\Upsilon_{\ell}\operatorname*{vec}(\tilde{\alpha}_{\ell}^{\prime}),$
where
$\Upsilon_{\ell}\equiv\Big{(}\operatorname*{vec}(\Xi_{i_{1}^{*},j_{1}^{*},h_{1},\ell}),\dots,\operatorname*{vec}(\Xi_{i_{k}^{*},j_{k}^{*},h_{k},\ell})\Big{)}^{\prime}\in\mathbb{R}^{k\times m^{2}}.$
Hence,
$\max_{\alpha(L)\colon\|\alpha(L)\|\leq M}\|R\operatorname{aBias}(\hat{\boldsymbol{\delta}})\|^{2}=\max_{\{\tilde{\alpha}_{\ell}\}_{\ell=1}^{\infty}\colon\sum_{\ell=1}^{\infty}\|\tilde{\alpha}_{\ell}^{\prime}\|^{2}\leq M^{2}}\left\|\sum_{\ell=1}^{\infty}R\Upsilon_{\ell}\operatorname*{vec}(\tilde{\alpha}_{\ell}^{\prime})\right\|^{2}.$
LABEL:thm:bias_general_aux1 shows that the final expression above equals $M^{2}\lambda_{\max}(\sum_{\ell=1}^{\infty}R\Upsilon_{\ell}\Upsilon_{\ell}^{\prime}R^{\prime})$ (the lemma only explicitly considers the case $M=1$, but the general case then follows from the homogeneity of degree 1 of the norm). Finally, Lemma B.1 below shows that $\sum_{\ell=1}^{\infty}\Upsilon_{\ell}\Upsilon_{\ell}^{\prime}=\operatorname{aVar}(\hat{\boldsymbol{\beta}})-\operatorname{aVar}(\hat{\boldsymbol{\delta}})$. This completes the proof of the proposition. The proof of LABEL:thm:bias_general_aux1 shows that the maximum above is achieved when $\operatorname*{vec}(\tilde{\alpha}_{\ell}^{\prime})\propto\Upsilon_{\ell}^{\prime}v$ (with the constant of proportionality being independent of $\ell$ and chosen to satisfy the norm constraint), where $v$ is the eigenvector corresponding to the largest eigenvalue of $R[\operatorname{aVar}(\hat{\boldsymbol{\beta}})-\operatorname{aVar}(\hat{\boldsymbol{\delta}})]R^{\prime}$. In the univariate case $k=1$, this reduces to expression (4.1) in Section 4.2. ∎
Lemma B.1.
Under the assumptions of Proposition 4.2, and using the notation in the proof of that proposition, we have
$\sum_{\ell=1}^{\infty}\Upsilon_{\ell}\Upsilon_{\ell}^{\prime}=\operatorname{aVar}(\hat{\boldsymbol{\beta}})-\operatorname{aVar}(\hat{\boldsymbol{\delta}}).$
Proof.
By definition of $\Upsilon_{\ell}$, it suffices to show that, for any indices $a,b\in\{1,\dots,k\}$,
$\sum_{\ell=1}^{\infty}\operatorname*{vec}(\Xi_{i_{a}^{*},j_{a}^{*},h_{a},\ell})^{\prime}\operatorname*{vec}(\Xi_{i_{b}^{*},j_{b}^{*},h_{b},\ell})=\operatorname{aCov}(\hat{\beta}_{i_{a}^{*},j_{a}^{*},h_{a}},\hat{\beta}_{i_{b}^{*},j_{b}^{*},h_{b}})-\operatorname{aCov}(\hat{\delta}_{i_{a}^{*},j_{a}^{*},h_{a}},\hat{\delta}_{i_{b}^{*},j_{b}^{*},h_{b}}).$ (B.2)
Multiplying out terms, we find that the left-hand side above equals
$\sum_{\ell=1}^{\infty}\operatorname*{trace}(\Xi_{i_{a}^{*},j_{a}^{*},h_{a},\ell}^{\prime}\Xi_{i_{b}^{*},j_{b}^{*},h_{b},\ell})$ $=\sum_{\ell=1}^{\infty}\operatorname*{trace}\left(A^{\ell-1}\Sigma(A^{\prime})^{\ell-1}S^{-1}\Psi_{i_{b}^{*},j_{b}^{*},h_{b}}\Sigma\Psi_{i_{a}^{*},j_{a}^{*},h_{a}}^{\prime}S^{-1}\right)$
$\qquad-\sum_{\ell=1}^{h_{b}}\operatorname*{trace}\left(A^{\ell-1}H_{\bullet,j_{b}^{*}}e_{i_{b}^{*},n}^{\prime}A^{h_{b}-\ell}\Sigma\Psi_{i_{a}^{*},j_{a}^{*},h_{a}}^{\prime}S^{-1}\right)$
$\qquad-\sum_{\ell=1}^{h_{a}}\operatorname*{trace}\left(A^{\ell-1}H_{\bullet,j_{a}^{*}}e_{i_{a}^{*},n}^{\prime}A^{h_{a}-\ell}\Sigma\Psi_{i_{b}^{*},j_{b}^{*},h_{b}}^{\prime}S^{-1}\right)$
$\qquad+\mathbbm{1}(j_{a}^{*}=j_{b}^{*})\sigma_{j_{a}^{*}}^{-2}\sum_{\ell=1}^{\min\{h_{a},h_{b}\}}e_{i_{b}^{*},n}^{\prime}A^{h_{b}-\ell}\Sigma(A^{\prime})^{h_{a}-\ell}e_{i_{a}^{*},n}.$
We now evaluate each of the four terms on the right-hand side above. The first term equals
$\operatorname*{trace}\Bigg{(}\underbrace{\sum_{\ell=1}^{\infty}A^{\ell-1}\Sigma(A^{\prime})^{\ell-1}}_{=S}S^{-1}\Psi_{i_{b}^{*},j_{b}^{*},h_{b}}\Sigma\Psi_{i_{a}^{*},j_{a}^{*},h_{a}}^{\prime}S^{-1}\Bigg{)}=\operatorname*{trace}\left(\Psi_{i_{b}^{*},j_{b}^{*},h_{b}}\Sigma\Psi_{i_{a}^{*},j_{a}^{*},h_{a}}^{\prime}S^{-1}\right).$
The second term (in the earlier display) equals
$-\operatorname*{trace}\Bigg{(}\underbrace{\sum_{\ell=1}^{h_{b}}A^{\ell-1}H_{\bullet,j_{b}^{*}}e_{i_{b}^{*},n}^{\prime}A^{h_{b}-\ell}}_{=\sum_{\ell=1}^{h_{b}}A^{h_{b}-\ell}H_{\bullet,j_{b}^{*}}e_{i_{b}^{*},n}^{\prime}A^{\ell-1}=\Psi_{i_{b}^{*},j_{b}^{*},h_{b}}}\Sigma\Psi_{i_{a}^{*},j_{a}^{*},h_{a}}^{\prime}S^{-1}\Bigg{)}=-\operatorname*{trace}\left(\Psi_{i_{b}^{*},j_{b}^{*},h_{b}}\Sigma\Psi_{i_{a}^{*},j_{a}^{*},h_{a}}^{\prime}S^{-1}\right),$
and the third term (in the earlier display) also equals this quantity by a symmetric calculation. In conclusion, we have shown
$\sum_{\ell=1}^{\infty}\operatorname*{trace}(\Xi_{i_{a}^{*},j_{a}^{*},h_{a},\ell}^{\prime}\Xi_{i_{b}^{*},j_{b}^{*},h_{b},\ell})$
$=\mathbbm{1}(j_{a}^{*}=j_{b}^{*})\sigma_{j_{a}^{*}}^{-2}\sum_{\ell=1}^{\min\{h_{a},h_{b}\}}e_{i_{b}^{*},n}^{\prime}A^{h_{b}-\ell}\Sigma(A^{\prime})^{h_{a}-\ell}e_{i_{a}^{*},n}-\operatorname*{trace}\left(\Psi_{i_{b}^{*},j_{b}^{*},h_{b}}\Sigma\Psi_{i_{a}^{*},j_{a}^{*},h_{a}}^{\prime}S^{-1}\right).$
The desired result (B.2) now follows from Corollary A.2. ∎
B.4 Proof of Corollary 3.2
Use the notation $E^{*}(z\mid w)=\operatorname*{Cov}(z,w)\operatorname*{Var}(w)^{-1}w$ for the mean-square projection of $z$ on $w$. Then
@@
where the last equality uses the assumption $\varepsilon_{j^{*},t-\ell}\in\operatorname{span}(\tilde{\check{y}}_{t-1},\dots,\tilde{\check{y}}_{t-p})$ for $\ell=1,\dots,h$. Thus,
@@
It now follows as in the proof of Corollary A.2 that $\operatorname{aVar}(\hat{\beta}_{h})=\operatorname{aVar}(\hat{\delta}_{h})$. Then Proposition 4.1 implies that $\operatorname{aBias}(\hat{\delta}_{h})=0$. ∎
B.5 Proof of Corollary 4.1
By Proposition 4.1,
$\sup_{\alpha(L)\colon\|\alpha(L)\|\leq M}\operatorname{aBias}(\hat{\delta}_{h};\alpha(L))^{2}=M^{2}\{\operatorname{aVar}(\hat{\beta}_{h})-\operatorname{aVar}(\hat{\delta}_{h})\}.$
Thus,
$\sup_{\alpha(L)\colon\|\alpha(L)\|\leq M}\operatorname{aMSE}(\hat{\delta}_{h};\alpha(L))-\operatorname{aMSE}(\hat{\beta}_{h})$
$=M^{2}\{\operatorname{aVar}(\hat{\beta}_{h})-\operatorname{aVar}(\hat{\delta}_{h})\}+\operatorname{aVar}(\hat{\delta}_{h})-\operatorname{aVar}(\hat{\beta}_{h})$
$=(M^{2}-1)\{\operatorname{aVar}(\hat{\beta}_{h})-\operatorname{aVar}(\hat{\delta}_{h})\}.\qed$
B.6 Proof of Corollary 4.2
Write $\hat{\theta}_{h}(\omega)=\hat{\delta}_{h}+\omega(\hat{\beta}_{h}-\hat{\delta}_{h})$. By Corollary A.2, the two terms are asymptotically independent of each other, and the second term has asymptotic variance $\omega^{2}\{\operatorname{aVar}(\hat{\beta}_{h})-\operatorname{aVar}(\hat{\delta}_{h})\}$. Hence,
@@
By Proposition 4.1, the supremum of the above expression over $\alpha(L)$ satisfying $\|\alpha(L)\|\leq M$ equals

### v3

@@ 6 Conclusion
    Researchers who use LP should control for those lags of the data that are strong predictors of the outcome or impulse variables. This is important not only when the shock is recursively identified, but even if the researcher directly observes a near-perfect proxy for the shock of interest. However, unlike for VAR inference, it is not necessary to get the lag length or set of control variables exactly right to achieve correct coverage. To select the number of lags to control for in the LP, we recommend running an auxiliary VAR in all variables used in the analysis and selecting the lag length to minimize the AIC; the auxiliary VAR is only used as a device to select the lag length and is otherwise discarded.
@@
    With the moderate lag lengths typical in current applied practice, VAR confidence intervals will only have accurate coverage at short horizons, and only because they are approximately equivalent with LP intervals at these horizons. If, at some horizons of interest, an estimated VAR yields confidence intervals that are substantially narrower than the corresponding LP intervals, we recommend increasing the VAR lag length until that is no longer the case, to guarantee robust confidence interval coverage. Conventional tests of correct VAR specification do not suffice to guard against coverage distortions.
Is there a way forward for VAR inference, beyond just including a large number of lags? We showed how to construct a VAR confidence interval with a bias-aware critical value that robustly controls coverage, but found that it will typically lead to wider confidence intervals than LP. Another option would be to estimate VARMA models rather than pure VARs, though this would be computationally expensive, and the bias-variance trade-off relative to LPs is unclear. In principle, VAR procedures may work better under additional restrictions on the misspecification, such as shape restrictions on the impulse response functions. [^11] Given any convex parameter space for the misspecification MA polynomial $\alpha(L)$, the worst-case bias of the VAR estimator (see Proposition 3.2) can be computed using convex programming. However, it appears that detailed application-specific restrictions would be required to generate a negligible worst-case bias, since we have shown that the least favorable misspecification in our baseline analysis cannot generally be ruled out based on economic theory. Rather than restricting the parameter space, future research could instead investigate weakening the coverage requirement, e.g., only requiring a certain coverage probability on average over a set of horizons (Armstrong, Kolesár, and Plagborg-Møller, 2022), or by changing the target for inference from the true impulse response function to a smooth projection of this function (Genovese and Wasserman, 2008). Finally, a subjectivist Bayesian VAR modeler need only worry about our negative results if their prior on potential misspecification attaches significant weight to MA processes that imply large VAR biases.
@@ A.1 Least favorable misspecification
Figure˜A.1 plots some numerical examples of the least favorable MA polynomial $\alpha^{\dagger}(L;h,M)=\sum_{\ell=1}^{\infty}\alpha^{\dagger}_{\ell,h,M}L^{\ell}$ discussed in Section˜4.2. We focus here on a univariate local-to-AR(1) model $y_{t}=\rho y_{t-1}+[1+T^{-1/2}\alpha(L)]\varepsilon_{t}$, though unreported numerical experiments suggest that the qualitative features mentioned below also apply to multivariate models. Recall that the least favorable MA coefficients depend on the horizon $h$ of interest, while $M$ only influences the overall scale of the coefficients, and not their shape as a function of $\ell$. The figure shows that the shape of the coefficients either takes the form of a hump or of a single zig-zag pattern, with the largest absolute value of the coefficients generally occurring at $\ell=h$. Notice that we can flip the signs of all coefficients without changing the absolute value of the bias.
@@ A.2 More efficient bias-aware confidence interval
Generalizing the bias-aware VAR confidence interval in Section˜4.3, consider a bias-aware confidence interval that is centered at the model averaging estimator $\hat{\theta}_{h}(\omega)=\omega\hat{\beta}_{h}+(1-\omega)\hat{\delta}_{h}$ from Corollary˜4.2:
@@
Impose Assumption˜2.1, $\zeta=1/2$, and $\operatorname{aVar}(\hat{\delta}_{h})>0$. Then, for any $\omega\in[0,1]$,
@@
The result follows from Propositions˜3.1, 3.2 and 4.1, Corollary˜A.2, and the same calculations as in the proof of Corollary˜4.2. ∎
@@
Figure˜A.3 shows this optimal weight as a function of $M$ and the relative asymptotic standard deviation of the VAR and LP estimators, while Figure˜A.3 shows the length of the resulting optimal bias-aware confidence interval relative to the length of the conventional LP interval. We see that, for $M\geq 2$, there is little gain from reporting the optimal bias-aware interval rather than the LP interval, regardless of the relative precision of VAR and LP. An additional observation is that, for $M\geq 1.5$, the length-optimal $\omega^{*}$ is numerically close to the MSE-optimal weight $M^{2}/(1+M^{2})$ derived in Corollary˜4.2.
Figure A.2: Length-optimal weight on LP in bias-aware confidence interval. Significance level $a=10\%$. Horizontal axis: relative asymptotic standard deviation of VAR vs. LP. Different lines: different bounds $M$ on $\|\alpha(L)\|$.
Figure A.3: Relative length of optimal bias-aware confidence interval vs. conventional LP interval. Significance level $a=10\%$. Horizontal axis: relative asymptotic standard deviation of VAR vs. LP. Different lines: different bounds $M$ on $\|\alpha(L)\|$.
@@ A.3 Covariance structure of LP and VAR estimators
The following result provides the asymptotic variance-covariance matrix of the LP and VAR estimators in the general multi-dimensional set-up of Section˜4.4. Define $\Psi_{i^{*},j^{*},h}$ as in Proposition˜3.2, but making the dependence on $(i^{*},j^{*})$ explicit in the notation.
@@
Impose Assumption˜2.1, with part (iii) holding for all shock indices $j_{1}^{*},\dots,j_{k}^{*}$. Then for any $a,b\in\{1,\dots,k\}$,
@@
and the $\operatorname{aCov}$” notation refers to elements of the asymptotic variance-covariance matrix in LABEL:eqn:asy_normal_joint in LABEL:app:joint. In particular, $\operatorname{aCov}(\hat{\beta}_{i_{a}^{*},j_{a}^{*},h_{a}}-\hat{\delta}_{i_{a}^{*},j_{a}^{*},h_{a}},\hat{\delta}_{i_{b}^{*},j_{b}^{*},h_{b}})=0$.
@@ Appendix B Proofs
Lemmas whose name begins with “E” can be found in LABEL:app:proof_details.
B.1 Proof of Proposition˜3.1
@@
$y_{i^{*},t+h}=\theta_{h,T}\varepsilon_{j^{*},t}+\underline{B}_{h,y}^{\prime}\underline{y}_{j^{*},t}+B_{h,y}^{\prime}y_{t-1}+\xi_{i^{*},h,t}+T^{-\zeta}\Theta_{h}(L)\varepsilon_{t},$ (B.1)
where the expressions for the coefficient matrices and the $1\times n$ two-sided lag polynomial $\Theta_{h}(L)=\sum_{\ell=-\infty}^{\infty}\Theta_{h,\ell}L^{\ell}$ are given in LABEL:thm:lp_coefficients_aux.
Let $\hat{x}_{h,t}$ be the residual in a regression of $y_{j^{*},t}$ on $\underline{y}_{j^{*},t}$ and $y_{t-1}$, using data points $1,2,\dots,T-h$. By definition, $\hat{x}_{h,t}$ is in-sample orthogonal to $\underline{y}_{j^{*},t}$ and $y_{t-1}$. Hence,
@@
$=\theta_{h,T}+\frac{T^{-1}\sum_{t=1}^{T-h}(y_{i^{*},t+h}-\theta_{h,T}\varepsilon_{j^{*},t}-\underline{B}_{h,y}^{\prime}\underline{y}_{j^{*},t}-B_{h,y}^{\prime}y_{t-1})\hat{x}_{h,t}+O_{p}(T^{-2\zeta})+o_{p}(T^{-1/2})}{\sigma_{j^{*}}^{2}+o_{p}(1)}$
by LABEL:thm:x(LABEL:itm:thm_x_iii)
@@
$=\theta_{h,T}+\frac{T^{-1}\sum_{t=1}^{T-h}(\xi_{i^{*},h,t}+T^{-\zeta}\Theta_{h}(L)\varepsilon_{t})\varepsilon_{j^{*},t}+O_{p}(T^{-2\zeta})+o_{p}(T^{-1/2})}{\sigma_{j^{*}}^{2}+o_{p}(1)}$
by LABEL:thm:x(LABEL:itm:thm_x_iv) and (LABEL:itm:thm_x_v)
@@
and the result follows. ∎
B.2 Proof of Proposition˜3.2
@@
LABEL:thm:A shows that $\hat{A}-A=O_{p}(T^{-\zeta}+T^{-1/2})$. By Magnus and Neudecker (2007, Table 7, p. 208),
@@
so
@@
where $S$ was defined in Assumption˜2.1. LABEL:thm:nu shows that
$\hat{\nu}-H_{\bullet,j^{*}}=\frac{1}{\sigma_{j^{*}}^{2}}T^{-1}\sum_{t=1}^{T}\xi_{0,t}\varepsilon_{j^{*},t}+o_{p}(T^{-\zeta}+T^{-1/2}).$
@@
B.3 Proof of Corollary˜3.2
Use the notation $E^{*}(z\mid w)=\operatorname*{Cov}(z,w)\operatorname*{Var}(w)^{-1}w$ for mean-square projection. Then
@@
where the last equality uses $\varepsilon_{j^{*},t-\ell}\in\operatorname{span}(\tilde{\check{y}}_{t-1},\dots,\tilde{\check{y}}_{t-p})$ for $\ell=1,\dots,h$. Thus,
@@
It now follows as in the proof of Corollary˜A.2 that $\operatorname{aVar}(\hat{\beta}_{h})=\operatorname{aVar}(\hat{\delta}_{h})$. Then Proposition˜4.1 implies that $\operatorname{aBias}(\hat{\delta}_{h})=0$. ∎
B.4 Proof of Corollary˜4.1
By Proposition˜4.1, $\sup_{\alpha(L)\colon\|\alpha(L)\|\leq M}\operatorname{aBias}(\hat{\delta}_{h};\alpha(L))^{2}=M^{2}\{\operatorname{aVar}(\hat{\beta}_{h})-\operatorname{aVar}(\hat{\delta}_{h})\}$. The result follows. ∎
B.5 Proof of Corollary˜4.2
Write $\hat{\theta}_{h}(\omega)=\hat{\delta}_{h}+\omega(\hat{\beta}_{h}-\hat{\delta}_{h})$. By Corollary˜A.2, the two terms are asymptotically independent of each other, and the second term has asymptotic variance $\omega^{2}\{\operatorname{aVar}(\hat{\beta}_{h})-\operatorname{aVar}(\hat{\delta}_{h})\}$. Hence,
@@
By Proposition˜4.1, the supremum of the above expression over $\alpha(L)$ satisfying $\|\alpha(L)\|\leq M$ equals